### v1

# An Empirical Evaluation of Pre-trained Large Language Models for Repairing Declarative Formal Specifications

Abstract.
Automatic Program Repair (APR) has garnered significant attention as a practical research domain focused on automatically fixing bugs in programs. While existing APR techniques primarily target imperative programming languages like C and Java, there is a growing need for effective solutions applicable to declarative software specification languages. This paper presents a systematic investigation into the capacity of Large Language Models (LLMs) for repairing declarative specifications in Alloy, a declarative formal language used for software specification. We propose a novel repair pipeline that integrates a dual-agent LLM framework, comprising a Repair Agent and a Prompt Agent. Through extensive empirical evaluation, we compare the effectiveness of LLM-based repair with state-of-the-art Alloy APR techniques on a comprehensive set of benchmarks. Our study reveals that LLMs, particularly GPT-4 variants, outperform existing techniques in terms of repair efficacy, albeit with a marginal increase in runtime and token usage. This research contributes to advancing the field of automatic repair for declarative specifications and highlights the promising potential of LLMs in this domain.
Declarative specification, Automatic Program Repair, formal methods, Alloy language, LLMs
copyright: none
1. Introduction
Declarative specification languages have become instrumental in addressing a multitude of software engineering challenges. Among these languages, the Alloy specification language (Jackson, 2006) has emerged as a powerful tool, leveraging relational algebra and first-order logic to tackle a diverse array of tasks within the software engineering domain. Its application spans across software verification (Bagheri et al., [n. d.]), the security analysis of cutting-edge platforms such as the IoT and Android systems (Bagheri et al., 2016, 2021; Alhanahnah et al., [n. d.]), and the creation of test cases (Khurshid and Marinov, 2004; Mirzaei et al., 2016). The integration of Alloy with the Alloy Analyzer has facilitated automated property verification, simplifying the process of checking whether specifications adhere to desired properties and seamlessly integrating it into the Alloy environment. Despite these advancements, Alloy users, akin to developers in imperative languages, encounter challenges in debugging and correcting subtle bugs that may arise during specification writing, particularly for complex systems.
While the Alloy Analyzer aids in automatic property verification and counterexample generation, debugging and rectifying issues in Alloy specifications remain laborious and manual tasks. Unlike the rich literature and techniques available for automatic program repair (APR) in imperative languages, the landscape for APR in declarative languages is relatively sparse. Early approaches like ARepair (Wang et al., 2019) primarily rely on failing test cases to identify and rectify bugs, assuming the presence of tests for verification. However, this paradigm does not align well with Alloy’s assertion-based specification approach, where users articulate expectations using assertions rather than tests. Existing APR techniques like ARepair (Wang et al., 2019) may succumb to overfitting issues when test coverage is insufficient, compromising the correctness of generated repairs. Despite efforts to mitigate these challenges with techniques like BeAFix (Brida et al., 2021), which leverages assertions as correctness oracles, there is room for improvement in terms of efficiency and effectiveness.
The emergence of Large Language Models (LLMs) has revolutionized various domains, including natural language processing and code generation. These pre-trained models, such as GPT-3.5-Turbo and GPT-4, exhibit remarkable accuracy in predicting natural language and generating code. Inspired by these advancements and the persistent challenges in repairing formal specifications, we embark on a systematic empirical study to explore the potential of utilizing LLMs for repairing faulty Alloy specifications. Our research aims to address key questions regarding the effectiveness, performance, adaptive prompting, failure characteristics, and repair costs associated with employing LLMs in Alloy specification repair.
This paper presents a methodical exploration and the results of our comprehensive evaluation of LLMs’ efficacy in repairing faulty Alloy specifications, juxtaposed against state-of-the-art Alloy APR techniques. We introduce a novel repair pipeline integrating a dual-agent LLM framework comprising a Repair Agent and a Prompt Agent  [^1] Our implementation and artifacts are available at: https://github.com/Mohannadcse/AlloySpecRepair. This iterative strategy centralizes around repairing bugs in defective specifications and generating specialized prompts to guide the repair process.
We conduct an extensive evaluation of LLM effectiveness in repairing flawed Alloy specifications, benchmarking it against several state-of-the-art Alloy APR techniques, including ARepair(Wang et al., 2019), ICEBAR (Gutiérrez Brida et al., 2023), BeAFix (Brida et al., 2021), and ATR (Zheng et al., 2022), shedding light on their capabilities, benefits, and potential limitations. Our evaluation encompassed two comprehensive sets of benchmarks, comprising 1,974 defective specifications sourced from ARepair (Wang et al., 2019) and Alloy4Fun (Macedo et al., 2021), developed by external research groups. The repair pipeline facilitates the rectification of each defective specification under six distinct scenarios, with up to six iterations per scenario. Our findings contribute to advancing the field of automatic repair for declarative specifications and provide insights into leveraging LLMs for effective Alloy specification repair.
2. Background and Motivation
This section provides an overview of multi-agent large language models followed by an illustrative example of a faulty Alloy specification to motivate the research and help the reader follow the discussions that ensue.
2.1. Multi-Agent Large Language Models
2.1.1. Large Language Models
Emerging pre-trained Large Language Models (LLMs) have demonstrated impressive performance on natural language tasks, such as text generation (Brown et al., 2020; Chowdhery et al., 2022; Vaswani et al., 2017) and conversations (Thoppilan et al., 2022; OpenAI, 2023). LLMs have also been proven effective in translating natural language specifications and instructions into executable code (Fan et al., 2023; Jain et al., 2022; Che, 2021). These models have been trained on extensive corpora and possess the ability to execute specific tasks without the need for additional training or hard coding (Bubeck et al., 2023). They are invoked and controlled simply by providing a natural language prompt. The degree to which LLMs understand tasks depends largely on the prompting techniques used to convey user-provided instructions. These prompts are categorized into two frameworks: “zero-shot” learning and “few-shot” learning (Brown et al., 2020). Within the “zero-shot” learning framework, the prompt includes a description of the problem and the instructions for its resolution, aiming that LLMs can tackle previously unseen tasks. Conversely, in the context of “few-shot” learning, LLMs are provided with examples, supplementing the guidance offered in the “zero-shot” prompt.
2.1.2. Multi-Agent LLMs.
With recent advancements in LLMs, developers are embracing the idea of creating autonomous agents that can solve various tasks and interact with environments, humans, and other agents using natural language interfaces (Zhou et al., 2023a). These agents provide several features including planning, memory, tool usage, and multi-agent communication. Therefore, LLM-based autonomous agents have gained tremendous interest in industry and academia (Wang et al., 2023). Several frameworks have been developed to support harnessing multi-agent LLM applications. AGENTS (Zhou et al., 2023b) is a unified framework and open-source library for language agents. AGENTS aims to facilitate developers to build applications with language agents. AutoGPT (Aut, [n. d.]) a multi-agent framework for LLMs, is designed to support autonomous applications of LLMs. LangChain (Chase, 2022) supports end-to-end language agents that can automatically solve tasks specified in natural language. It also facilitates the connection between LLM agents and external tools. Langroid (lan, 2024) is another framework that supports the development of multi-agent LLMs. It offers the capability to manage the history of messages, thus controlling the context window of LLMs. This is a crucial functionality for LLM apps operating iteratively (i.e., Self-Refine (Madaan et al., 2023) and Reflexion (Shinn et al., 2024)). Moreover, Langroid enables seamless integration with a variety of LLMs.
2.1.3. tools and function calling
Multi-agent frameworks and recent versions of OpenAI’s LLM have introduced a feature known as function calling [^2] https://platform.openai.com/docs/guides/text-generation/function-calling. This feature enables users to provide function descriptions to the LLM. In turn, the LLM responds with a structured response (i.e., JSON data containing the requisite arguments for invoking any available functions). These functions serve as action executors and provide the option to supply APIs that the LLM can query to obtain essential information for responding to users. For example, when a user inquires about the current weather in a specific city and provides the LLM with a weather API call description, the LLM can augment its response with information retrieved from the API. This facilitates a seamless and efficient interaction between the user and the LLM and provides rich responses to avoid hallucinations.
2.2. Motivating Example
Alloy, a formal modeling language, provides a comprehensible syntax inspired by object-oriented notations and is grounded on first-order relational logic (All, [n. d.]). Within an Alloy specification, three primary components shape its structure: data types, constraints expressed through formulas, and commands to initiate the analyzer.
The language relies on sig to define fundamental data types, which can be organized using fields to capture relationships between signatures. Additionally, Alloy employs fact to introduce constraints that persist in every instance of the specification. These constraints restrict the model space, ensuring its consistency. Further structuring of formulas is achieved through pred and fun, which are named parameterized Alloy formulas, and assert encapsulates the properties intended for validation.
Commands such as check and run activate the automated analyzer. Check verifies assertions, while run executes predicates, aiming to identify model instances that satisfy specified conditions. Alloy’s expressiveness stems from its use of relational logic, a first-order logic extended with relational operators. These include all, some, one, and lone quantifiers, along with operators like relational join and transitive closure.
To illustrate the Alloy language and motivate the research, we examine the specification of the River Crossing puzzle, a classic problem of safely transporting a farmer, a fox, a chicken, and a sack of grain across a river. Listing 1 presents an abbreviated Alloy specification snippet of the farmer model from ARepair benchmark (Wang et al., 2019). In the River Crossing puzzle, the goal is to transport all objects from one side of the river to the other safely. However, there are constraints to consider: the farmer can only take one item with them in the boat, and if the farmer leaves the fox with the chicken, or the chicken with the grain, without supervision, a dire consequence will occur. This puzzle’s Alloy specification defines objects like the Farmer, Fox, Chicken, and Grain, along with their attributes and constraints.
However, the model contains a flaw in the crossRiver predicate, where the logic mistakenly assumes the occurrence of eating actions on the wrong side of the river. This oversight allows scenarios where the fox can eat the chicken without the farmer’s supervision, violating the puzzle’s rules. Interestingly, state-of-the-art automated repair techniques have struggled to address this particular bug. To repair this model, the erroneous logic within the crossRiver predicate must be corrected. However, this task presents a significant challenge due to the complexity of the puzzle’s logic and interactions between its components. Rectifying these logical errors requires a deep understanding of the problem domain and intricate relationships within the specification.
The objective of our study is to assess the effectiveness of repair methods for such buggy specifications using LLMs compared to state-of-the-art Alloy repair techniques. LLMs offer potential in uncovering complex violations and evaluating repair methods by reasoning about broader system contexts and intended behaviors. In the following sections, we explore the ability of LLMs to address such challenges and the specific conditions under which they can do so.
1abstract sig Object { eats: set Object }
2one sig Farmer, Fox, Chicken, Grain extends Object {}
3fact eating { eats = Fox->Chicken + Chicken->Grain }
4sig State {
5 near: set Object,
6 far: set Object
7}
8fact initialState {
9 let s0 = ord/first |
10 s0.near = Object && no s0.far
11}
12// Bug found in original model crossRiver
13pred crossRiver [from, from’, to, to’: set Object] {
14 // Fix: replace "from’ = from - Farmer && to’ = to - to.eats + Farmer" with "from’ = from - Farmer - from’.eats && to’ = to + Farmer".
15 (from = from - Farmer && to = to - to.eats + Farmer)
16 // Fix: replace "from’ = from - Farmer - item && to’ = to - to.eats + Farmer + item" with "from’ = from - Farmer - item - from’.eats && to’ = to + Farmer + item".
17 || (some item: from - Farmer {
18 from = from - Farmer - item
19 to = to - to.eats + Farmer + item })
20}
21fact stateTransition {
22 all s: State, s’: ord/next[s] {
23 Farmer in s.near =>
24 crossRiver[s.near, s’.near, s.far, s’.far] else
25 crossRiver[s.far, s’.far, s.near, s’.near]
26 }
27}
28pred solvePuzzle { ord/last.far = Object }
29run solvePuzzle for 8 State expect 1
30assert NoQuantumObjects {
31 no s : State | some x : Object | x in s.near and x in s.far
32}
33check NoQuantumObjects for 8 State expect 0
Listing 1: A snippet of the flawed farmer Alloy specification sourced from the ARepair benchmark [^4] https://github.com/guolong-zheng/atmprep/blob/master/benchmark/arepair/farmer1.als.
3. Methodology
This section outlines the repair pipeline, detailing its architecture and workflow. The pipeline is driven by the APR phases as outlined in (Zhang et al., 2023), including localization, repair, and verification. Our implementation of this pipeline utilizes the Langroid framework (lan, 2024).
Figure 1. LLM-assisted APR Pipeline.
3.1. Overall Workflow
In this section, we overview our repair pipeline specifically tailored for iterative self-refinement. Our approach revolves around the concept of iterative refinement, wherein the repair process for a defective Alloy model operates within a predetermined budget, defined by the number of iterations allocated. Previous research has indicated that this iterative prompting method yields superior outcomes compared to traditional single-step approaches, enhancing task efficiency by an average of 20
The workflow of our Automatic Program Repair (APR) pipeline, as depicted in Figure 1, comprises two primary components: the Repair Agent and the Prompt Agent. Each agent maintains its context, tools, and prompts, but they collaborate to adaptively refine the prompt for enhanced effectiveness, as elaborated below. Our pipeline supports various feedback levels, employing a zero-shot prompting approach (Paul et al., 2023), where tasks are performed without explicit examples, as detailed in Table 1.
Upon receiving defective Alloy specifications and the prompt, our APR pipeline dispatches these elements to the LLM, which generates a revised version aimed at rectifying the defect. Subsequently, the Repair Agent compares the proposed specifications with the buggy ones to evaluate LLM’s ability to localize faults and adhere to instructions. The correctness of the proposed specifications is then assessed by triggering the Alloy Analyzer tool, which runs the proposed specifications for validation. If the Alloy analyzer confirms the bug’s resolution, the repair process concludes. Conversely, if errors persist, the pipeline initiates another iteration, contingent upon the remaining repair budget. If the budget is depleted, the pipeline terminates operations.
Our adaptive prompt is designed to evolve and refine based on different settings (detailed in Section 3.3), ensuring continual improvement after each unsuccessful attempt.
The following sections provide a comprehensive overview of each component of the pipeline.
Table 1. Zero-shot prompts used by the repair pipeline
Agent Message Type Message content
Repair Agent Agent-Instruction You are an expert in repairing Alloy declarative specifications. You will be presented with Alloy ¡Faulty_SPECIFICATIONS¿. Your task is to FIX/REPAIR the ¡Faulty_SPECIFICATIONS¿. Use the tool ‘run_alloy_analyzer‘ to demonstrate and validate the ¡FIXED_SPECIFICATIONS¿. Wait for my feedback, which may include error messages or Alloy solver results. You will have 5 trials to fix the ¡Faulty_SPECIFICATIONS¿. **Adhere to the Following Rules**: - The ¡FIXED_SPECIFICATIONS¿ should be consistent (having instances) and all the assertions should be valid (no counterexample). - DO NOT REPEAT the ¡FIXED_SPECIFICATIONS¿ that I sent you. - DO NOT provide any commentary and always send me anything ONLY using the tool ‘run_alloy_analyzer‘. - The ¡FIXED_SPECIFICATIONS¿ MUST be different than the ¡Faulty_SPECIFICATIONS¿.
Tool-instruction ALL AVAILABLE TOOLS and THEIR JSON FORMAT INSTRUCTIONS: You have access to the following TOOLS to accomplish your task: TOOL: run_alloy_analyzer, PURPOSE: To show a ¡FIXED_SPECIFICATIONS¿ to the user. Use this tool whenever you want to SHOW or VALIDATE the ¡FIXED_SPECIFICATIONS¿. NEVER list out a ¡FIXED_SPECIFICATIONS¿ without using this tool. JSON FORMAT: ”type”: ”object”, ”properties”: ”request”: ”default”: ”run_alloy_analyzer”, ”type”: ”string”,”specification”: ”type”: ”string”, ”required”: [ ”specification”, ”request”]. When one of the above TOOLs is applicable, you must express your request as ”TOOL:” followed by the request in the above JSON format.
Prompt Agent Agent-Instruction You are Expert in Analyzing Alloy Analyzer reports. Can you describe concisely and precisely the modifications needed to fix the error in at most 2 sentences? Based on this report from Alloy Analyzer: {Alloy_report_msg} After running this Alloy Model is: {proposed_spec}
3.2. Pipeline Agents
As depicted in Figure 1, the pipeline comprises two agents. This design facilitates evaluating the recent deployment of LLM apps as multi-agent applications.
3.2.1. Repair Agent
This agent is the core component of the APR pipeline. It has access to tools/functions (i.e., “run_alloy_analyzer”), maintains its context (i.e., history of messages) over the iterations, and decides the process for handling subsequent iterations.
The agent’s prompt encompasses a set of messages combined as a system message. These messages are listed in Table 1.
  • Agent Instruction. This message provides a general guideline to the Repair Agent, commencing with the agent’s persona as an “expert in repairing Alloy specifications”. It proceeds to notify LLM of defects within the provided Alloy model, without disclosing the defects’ locations. The message finally describes the tools (i.e., “run_alloy_analyzer”) to which this agent has access. In conclusion, it enumerates a set of procedural steps to be executed by the LLM.
  • Tool Instruction. Offers instructions delineate the conditions and procedures for utilizing the tool “run_alloy_analyzer”. It further specifies the data required from the LLM to ensure the proper activation of this tool. This instruction guides the LLM to transmit a JSON formatted response, encapsulating two fields: (1) request, denoting the tool designation intended for use by the LLM, and (2) proposed_specification, representing the LLM’s suggested version for rectifying the identified bug.
3.2.2. Prompt Agent
This agent has its LLM settings and context and generates feedback based on the report generated by Alloy Analyzer and the proposed specification. The produced feedback guides the Repair Agent on the methods for rectifying the bug. The system message that provides the instructions to this agent is described in Table 1. Unlike the system message of the Repair Agent, the agent’s persona is described as “Expert in Analyzing Alloy Analyzer reports”. This agent does not share the same context or prompts with the Repair Agent. But the response of the Prompt Agent is appended to the prompt that will be used in the next iteration by the Repair Agent.
3.3. Feedback Messages
The initial prompt contains only Agent Instruction and Tool Instruction defined by the Repair Agent, as described in Table 1. The pipeline refines the prompts in response to the behavior of the proposed specification. This refinement process is contingent upon repetitive buggy specifications, errors that emerge after the execution of these specifications through the Alloy analyzer, or LLM failure to send a response based on the required JSON format.
To this end, Table 2 presents the feedback messages that are supported by the pipeline. These feedback messages will be appended to the initial prompt. Following is a description of these:
  • Tool fallback. This message informs LLM that the response received does not comply with the tool’s JSON format, consequently impeding the extraction of the proposed specification. To address this issue, we developed a parser atop the Langroid JSON parser, specifically to retrieve Alloy specifications from the response. Nonetheless, the parser cannot handle all scenarios, resulting in the dispatch of this feedback to the LLM, soliciting adherence to the mandated JSON format.
  • Repeated Spec. This message is triggered when the proposed spec is the same as the buggy specification. In this situation, this message tells the LLM not to repeat the buggy specification.
  • Alloy Analyzer Report. This message relays the feedback of the verification process using Alloy Analyzer. Table 2 presents three feedback levels.
In the following, we discuss the details of the various feedback levels supported by the pipeline.
Table 2. Feedback Messages used by the repair pipeline
Feedback Type Generated By Message Content
Tool-fallback Alloy Analyzer You must use the CORRECT format described in the tool ‘run_alloy_analyzer‘ to send me the fixed specifications. You either forgot to use it, or you used it with the WRONG format. Make sure all fields are filled out.
Repeated spec The proposed ¡FIXED_SPECIFICATIONS¿ is IDENTICAL to the Alloy ¡Faulty_SPECIFICATIONS¿ that I sent you. **DO NOT** send Alloy specifications that I sent you again. ALWAYS USE the tool ‘run_alloy_analyzer‘ to send me a new ¡FIXED_SPECIFICATIONS¿.
No-feedback The proposed specification DID NOT fix the bug.
Generic-feedback Below are the results from the Alloy Analyzer. Fix all Errors and Counterexamples before sending me the next ¡FIXED_SPECIFICATIONS¿.
Auto-feedback Prompt Agent The message content will be dynamically constructed by the agent based on the supplied information.
Alloy Analyzer Feedback Levels: The verification stage involves running the proposed specifications with the Alloy Analyzer. Subsequently, feedback conveying the Alloy Analyzer’s report is generated, available at different levels of detail. This process aids in assessing the effectiveness of various refinements made to the prompts.
In this study, we consider three feedback levels to refine the prompt as outlined in Table 2. These levels reflect errors, counterexamples, and instances identified after running the Alloy Analyzer in each iteration. They also mimic different repair scenarios, as described below:
  1. (1)
    No-feedback. In this setting, the Alloy Analyzer agent returns only a single response (i.e., The proposed specification DID NOT fix the bug.) to LLM without describing the details of the report generated by the Alloy Analyzer. This setting is designed to demonstrate the ability of LLMs to identify and fix errors in a faulty model solely based on the content of the buggy model itself, mirroring the function of a programming language compiler.
  2. (2)
    Generic-feedback. We use a template to send the Alloy Analyzer report to LLM. This report summarizes counterexamples, instances, and errors. This scenario represents a common situation where the developer shares a summary of the issue on a question-and-answer platform (such as Stack Overflow) to get help.
  3. (3)
    Auto-feedback. When this setting is enabled, the feedback response to LLM will be generated by another LLM agent (i.e., prompt agent). We forward the report generated by the Alloy Analyzer and the proposed Alloy specifications to the prompt agent, which generates a prompt to instruct the Repair Agent.
The different reports reflect the ability of LLMs to (1) locate bugs and (2) correctly perform the repair process. For example, Generic-feedback and Auto-feedback show the ability of LLM to perform its reasoning based on a human-created prompt versus an LLM-generated prompt, which has no access to the repair context. Also, the latter reflects a real scenario wherein the user cannot repair the buggy specification, then they consult experts or other forums to get assistance.
4. Experiment Design
In this study, we address the following research questions (RQs):
  • RQ1 (Effectiveness): How effective is the APR pipeline in repairing compared to the state of the art?
  • RQ2 (Performance): What is the repair performance when employing pre-trained LLMs?
  • RQ3 (Adaptive Prompting): What is the impact of applying the adaptive prompting technique?
  • RQ4 (Failure Characteristics): What are the characteristics of failures encountered during the repair process?
  • RQ5 (Repair Costs): What is the cost associated with using the APR pipeline with various LLMs?
4.1. Dataset
In our evaluation, we utilize two distinct benchmark suites: ARepair (Wang et al., 2019) and Alloy4Fun (Macedo et al., 2021). These benchmark suites have undergone extensive study and been developed by independent research groups, enabling a fair comparison of various techniques.
The benchmark datasets consist of specifications varying from tens to hundreds of lines, featuring real bugs authored by humans. The defects within the benchmarks span a diverse range, from straightforward issues that can be addressed by modifying a single operator to complex challenges necessitating the synthesis of novel expressions and the substitution of entire predicate bodies.
The ARepair benchmark (Wang et al., 2019) encompasses 38 faulty specifications extracted from a total of 12 Alloy problems. Within this benchmark, both single and multi-line errors are present, distributed across 28 faulty models with single-line errors and 10 faulty models with multi-line errors. The Alloy4Fun benchmark (Macedo et al., 2021) comprises a collection of 1,938 handwritten defective models sourced from student submissions across six different Alloy problems. All bugs within this benchmark are single-line bugs. Both benchmarks are accompanied by correct versions of the specifications, serving as ground truths for verifying the accuracy of the obtained results.
4.2. Pre-processing Benchmark Dataset
As mentioned earlier, the benchmark datasets include comments indicating bug locations and their corresponding fixes. To ensure a fair evaluation, we have removed these comments from the defective models. Specifically, we have eliminated all lines starting with the phrase ”Fix:” in the defective models. This line typically outlines the necessary changes required to resolve the bug. Models with multiple ”Fix:” comments are categorized as having multi-line bugs, whereas those with only one ”Fix:” comment are considered to have single-line bugs.
This process of removing comments is crucial as it prevents LLMs from being provided with explicit clues about the locations of bugs and their respective fixes. By doing so, we aim to accurately assess the repair capabilities of LLMs and observe their effectiveness in localizing bugs within the defective models.
4.3. Implementation
Our implementation of the APR pipeline was realized using Python 3.11, with the assistance of Langroid to facilitate the integration of multi-agent LLMs. This implementation encompasses several crucial functionalities. Firstly, it offers support for various LLMs, including local LLMs, ensuring flexibility in model selection. Secondly, it incorporates a message history control mechanism, essential for preventing context length limitations, particularly during iterative repair processes. Lastly, our implementation facilitates the creation and customization of tools, allowing for the definition of response formats and fields. Additionally, we developed a specialized parser to address issues caused by special situations that could impede the extraction of proposed specifications.
The experiments were carried out on a system equipped with a 2.3 GHz Quad-Core Intel Core i7 processor, 32 GB of RAM, and running macOS Sonoma. Furthermore, the system was configured with Oracle Java SE Development Kit 8u202 (64-bit version), ensuring compatibility and optimal performance throughout the execution of the APR pipeline.
4.4. Subject LLMs
To assess the performance of existing pre-trained LLMs, we have carefully chosen three representative models [^5] While exploring local models such as Llama-2 and Mistral, it became apparent that they lacked support for tools, making it challenging to process their responses, which often contained incomplete Alloy specifications.:
  • GPT-3.5-Turbo: This standard LLM, provided by OpenAI and utilized in ChatGPT, boasts a sophisticated architecture with 175 billion parameters, endowing it with extensive capabilities. Engineered to tackle a wide array of natural language processing (NLP) tasks, including text generation and completion, GPT-3.5-Turbo epitomizes advanced computational linguistics. Within the context of our study, we utilize the gpt-3.5-Turbo variant accessed via the OpenAI API for evaluative purposes.
  • GPT-4-32k: Representing the state-of-the-art in LLM models from OpenAI, this model contains over 1.76 trillion parameters, demonstrating significantly enhanced capabilities compared to GPT-3.5. For our study, we leverage the GPT-4-32k-0613 model version.
  • GPT-4-Turbo: Another iteration of the GPT-4 model, GPT-4-Turbo features an updated knowledge cutoff as of April 2023 and introduces a 128k context window. Moreover, it offers cost-effectiveness compared to GPT-4, alongside notable enhancements such as improved instruction following, JSON mode, and reproducible outputs (OpenAI, [n. d.]).
Table 3 provides more details about the characteristics of these pre-trained LLMs.
Table 3. Characteristics of Pre-trained LLMs
Model Version Cut-off
Context Window
Input Cost
per 1M tokens
Output Cost
per 1M tokens
GPT-3.5 Turbo 1106 Sep 2021 16,385 \$1 \$2
GPT-4-32k 0613 Sep 2021 32,768 \$60 \$120
GPT-4 Turbo 1106-preview Apr 2023 128k \$10 \$30
4.5. Subject SOTA Systems
We conducted a comparative analysis of repair performance against several state-of-the-art Alloy repair tools: ARepair (Wang et al., 2019), ICEBAR (Gutiérrez Brida et al., 2023), BeAFix (Brida et al., 2021), ATR (Zheng et al., 2022), and Hasan et al. (Hasan et al., 2023). Each tool employs a distinct development approach and tackles specification attributes differently.
ARepair (Wang et al., 2019) generates fixes for Alloy specifications that violate test cases. ICEBAR (Gutiérrez Brida et al., 2023) utilizes ARepair as a backend tool to repair faulty Alloy specifications based on a predefined set of Alloy tests. ATR (Zheng et al., 2022) adopts a template-based methodology to enhance the repair process, leveraging fault localization strategies to identify potentially erroneous Alloy expressions leading to assertion failures. BeAFix (Brida et al., 2021) relies on user input to identify faulty statements. It exhaustively explores all possible repair candidates up to a certain threshold through mutation and employs Alloy counterexamples to evaluate the feasibility of generalization. Fianlly, Hasan et al. (Hasan et al., 2023) employed ChatGPT for repairing faulty specifications across five scenarios. For our comparison, we exclude scenarios where ChatGPT is provided with “Fix:” comments, as this represents an unrealistic setting and does not align with our ”Benchmark Pre-processing” step (cf. Section 4.2).
4.6. Experimental Configuration and Settings
LLM Temperature: To balance between deterministic progression and iterative refinement, we set the temperature parameter of the Large Language Models (LLMs) to 0.2, allowing for a moderate level of randomness while maintaining some level of determinism throughout the repair process (Kong et al., 2024; Zhang et al., 2024).
Number of Iterations: Initially, we conducted a preliminary experiment using the ARepair benchmark framework with GPT-4-32k, allocating a budget of ten iterations. However, we observed diminishing returns after six iterations in most cases. Therefore, we opted for a six-iteration budget in the APR pipeline.
Metric for Comparing Repair Performance of LLMs: We employ the Correct@k metric (Liu et al., 2024) to evaluate the success rate of the techniques in repairing defective Alloy specifications. This metric quantifies the number of defects successfully repaired within a maximum of $k$ iterations, where $k$ is set to 6 in our study (denoted as Correct@6). The formula for this metric is described in Equation 1.
[…]
Experiment Settings and Baseline: Our evaluation encompasses a total of 1,974 defective models. We assess three distinct LLMs, as outlined in Table 3. The APR pipeline supports three levels of report granularity and permits multiple iterations, resulting in a total of 106,596 repair attempts (calculated as the product of 1,974 benchmarks, three LLMs, three feedback levels, and up to six iterations). To manage costs associated with GPT-4 models, we conducted two experiments for RQ1. The first experiment focuses on repairing Alloy4Fun and ARepair benchmarks using GPT-3.5, followed by a comparison with state-of-the-art (SoTA) tools. In the second experiment, we limit the repair task to the ARepair benchmark, utilizing GPT-4 and GPT-4-Turbo exclusively. Table 4 provides details of the different combinations of LLMs and feedback intensities used in the latter experiment.
Table 4. Settings corresponding to various combinations of LLMs and feedback levels
Setting No. LLM Feedback Level
Setting-1 GPT-4-32k No-Feedback
Setting-2 GPT-4-32k Generic-Feedback
Setting-3 GPT-4-32k Auto-Feedback
Setting-4 GPT-4-Turbo No-Feedback
Setting-5 GPT-4-Turbo Generic-Feedback
Setting-6 GPT-4-Turbo Auto-Feedback
5. Experimental Results
This section summarizes the data we collected, its interpretation, and our results.
5.1. Results for RQ1: Effectiveness
Table 5. Comparing against state-of-the-art Alloy repair techniques (ARepair (Wang et al., 2019), ICEBAR (Gutiérrez Brida et al., 2023), BeAFix (Brida et al., 2021), ATR (Zheng et al., 2022), and Hasan et al. (Hasan et al., 2023)) on the ARepair and Alloy4Fun benchmarks, using GPT-3.5-Turbo as our LLM baseline.
Total Arepair ICEBAR BeAFix ATR Hasan et al. No-Feedback Generic-Feedback Auto-Feedback
#specs #repairs #repairs #repairs #repairs #repairs #repairs #repairs #repairs
Alloy4Fun
1936 185 1051 981 1286 147 156 231 278
ARepair
38 9 21 24 22 4 4 6 18
Table 6. Comparing with state-of-the-art Alloy repair techniques (ARepair (Wang et al., 2019), ICEBAR (Gutiérrez Brida et al., 2023), BeAFix (Brida et al., 2021), ATR (Zheng et al., 2022), and Hasan et al. (Hasan et al., 2023)) on the ARepair benchmarks, using various combinations of LLMs and report levels as outlined in Table 4.
Model Total Arepair ICEBAR BeAFix ATR Hasan et al. Setting-1 Setting-2 Setting-3 Setting-4 Setting-5 Setting-6
#specs #repairs #repairs #repairs #repairs #repairs #repairs #repairs #repairs #repairs #repairs #repairs
ARepair addr 1 1 1 1 1 0 0 0 1 0 1 1
arr 2 2 2 2 1 0 0 0 0 1 0 1
balancedBST 3 1 2 1 1 1 1 1 2 1 1 1
bempl 1 0 1 0 1 0 0 0 0 1 1 1
cd 2 0 2 2 2 1 1 2 2 2 1 2
ctree 1 1 0 0 0 1 1 1 1 1 1 1
dll 4 0 3 3 2 4 4 4 4 2 2 4
farmer 1 0 0 0 0 1 1 1 1 1 1 1
fsm 2 2 2 1 2 1 1 1 2 0 0 1
grade 1 0 1 0 1 0 0 0 0 0 0 1
other 1 0 0 1 1 0 0 0 0 0 0 1
student 19 2 7 13 10 6 6 6 9 8 11 13
Summary 38 9 21 24 22 15 15 16 22 17 19 28
Figure 2. Venn diagrams illustrating the exclusive and overlapping cases of successful repairs by different repair methods, highlighting the complementary and unique capabilities of the APR pipeline compared to SoTA repair tools.
Table 5 provides a comparison using GPT-3.5-Turbo as the baseline. The column ”Total #specs” indicates the total number of defective specifications in each benchmark, while the subsequent columns show the number of correctly repaired specifications by each state-of-the-art Alloy repair tool. The last three columns display the repair outcomes based on different report levels employed in the pipeline. The performance of the baseline LLM (GPT-3.5-Turbo) is notably inferior compared to traditional state-of-the-art tools, consistent with prior findings by Hasan et al. (Hasan et al., 2023). However, the two-agent (auto-feedback) configuration demonstrated the most promising performance among the feedback levels across both benchmarks.
Table 6 presents comparison results for the ARepair benchmark using GPT-4-32k and GPT-4-Turbo. The ”Models” column categorizes model groups in the ARepair benchmark, while the last six columns illustrate the total count of accurately fixed specifications under different pipeline settings (summarized in Table 4). Setting-6 notably outperforms all state-of-the-art tools, while Setting-3 performs similarly to BeAFix and surpasses other traditional tools. It’s noteworthy that the specification farmer remained unresolved by all SoTA repair tools, but was successfully repaired in the APR pipeline across all settings. This suggests that GPT-4 family models can effectively address complex specifications, particularly those with multi-line bugs. Further, the results observed in Setting-3 and Setting-6 surpass those of other settings (in Table 4) and align with the ”Auto-Feedback” column in Table 5. This alignment underscores the efficacy of prompt construction using LLMs over generic prompts developed by humans, advocating for the use of multi-agent LLM setups.
To gain deeper insights into the complementarity of these methods, Venn diagrams in Figure 2 illustrate cases exclusively fixed by one tool but not by others, as well as the overlap between techniques (i.e., cases correctly repaired by multiple tools). Notably, the repair pipeline successfully addresses faulty models that other SoTA repair tools fail to resolve.
The metric Correct@6 is utilized to compare the performance of all LLMs across various feedback levels. Table 7 showcases the comparison results using the ARepair benchmark. While GPT-3.5-Turbo exhibited less effective repair capabilities, GPT-4-Turbo demonstrated superior performance across different feedback levels, attributable to its cut-off and context window parameters relative to its counterparts (see Table 3), alongside enhanced features integrated into this model (detailed in Section 4.4).
The results for RQ1 demonstrate that the APR pipeline, particularly with GPT-4-Turbo and multi-agent setups, outperforms traditional state-of-the-art Alloy repair tools, effectively addressing complex multi-line bugs and showcasing superior performance across various feedback levels.
Table 7. Comparison results presenting the repair performance of different LLMs across various feedback levels, i.e., No-Feedback (NF), Generic-Feedback (GF), Auto-Feedback (AF), measured by the Correct@6 metric.
GPT-3.5-Turbo GPT-4-32k GPT-4-Turbo
NF GF AF NF GF AF NF GF AF
10.5 15.8 47.4 39.5 42.1 57.9 44.7 50.0 57.9
5.2. Results for RQ2: Performance
This research question delves into the repair performance of various pre-trained Language Model Models (LLMs) concerning factors such as repair iteration budget, bug type, localization capabilities, and adherence to instructions. Our analysis focuses on two LLM variants: GPT-4-32k and GPT-4-Turbo, utilizing the results provided in Table 6 and the settings outlined in Table 4.
Repair Iteration Budget: The box-whisker plot in Figure 3(a) illustrates the iterations needed by different settings to rectify faulty specifications. Across all settings, the median number of iterations predominantly centers around 1.0, except for Setting 3 & 6, where Setting-3 increases to 2.5, indicating a demand for additional iterations. However, median values for Settings 1, 2, 4, and 5 remain at 1.0, suggesting that at least half of the issues are resolved within the first iteration.
(a) Refer to caption
(b) Refer to caption
(c) Refer to caption
Figure 3. (a) Iteration count distribution for repairing specifications across various settings. (b) Repair distribution in ARepair benchmark categorized by bug type: single-line or multi-line (higher values preferred). (c) The incident count for initial repair attempts mirroring the buggy specification (lower values preferred), derived from the ARepair benchmark.
Repairing Based on Bug Type: Figure 3(b) depicts the repair capabilities of GPT-4-32k and GPT-4-Turbo concerning bug types, namely single-line and multi-line bugs. The best performance for GPT-4-Turbo (auto-feedback, Setting-6) achieved the highest repair rate, fixing 16 single-line bugs and 12 multi-line bugs. In contrast, the best result for GPT-4-32k (auto-feedback, Setting-3) demonstrated superior performance in correcting single-line bugs (12 models) compared to multi-line bugs (10 models). Additionally, we examine the repair outcomes achieved by GPT-3.5-Turbo, showing repair capabilities for both single-line and multi-line bugs.
Following Instructions: Adherence to instructions is crucial in tasks like APR to ensure repair quality and control costs. However, Figure 3(c) reveals that both GPT-4 models sometimes repeat buggy models in their initial repair iterations, despite system instructions to avoid this. GPT-4-Turbo generally shows better compliance, particularly in Setting-6, where it exhibits the lowest number of repeated buggy specifications.
The comparison between the APR pipeline utilizing different feedback levels and state-of-the-art Alloy repair techniques, along with the performance of various LLMs, revealed that the auto-feedback approach exhibited the best performance across both benchmarks, surpassing traditional SoTA repair tools.
5.3. Results for RQ3: Adaptive Prompting
The APR pipeline facilitates three feedback levels—No-feedback, Generic-Feedback, and Auto-Feedback—as outlined in Section 2. This research question delves into the implications of these feedback levels on LLM performance, drawing insights from the following analysis:
Overall Repair Performance: Examination of Tables 5 and 6 reveals a notable correlation between the use of ”auto-feedback” and superior repair efficacy across all LLMs, including GPT-3.5-Turbo. Conversely, the absence of feedback, characterized by ”No-feedback,” resulted in the poorest repair performance, indicating that prompt updates driven by information from the Alloy Analyzer report significantly contribute to repair success.
Bug-type Repair: Across all LLMs, Settings 3 and 6 consistently demonstrate a higher ability to repair complex specifications, as illustrated in Figure 3(b). This suggests that adaptive prompting, particularly in Settings 3 and 6, enhances the pipeline’s capacity to address intricate bugs effectively.
Iteration Budget Analysis: Figure 3(a) showcases a slight increase in the requirement for additional iterations in Settings 3 and 6. While these settings exhibit superior repair capabilities, they may necessitate a marginally higher number of iterations to achieve optimal outcomes.
These observations underscore the advantage of leveraging LLMs to automatically generate prompts based on Alloy Analyzer reports, leading to enhanced repair performance compared to Generic-Feedback. However, this advantage may be accompanied by the need for additional iterations.
5.4. Results for RQ4: Characteristics of Failed Repairs
This research inquiry delves into the root causes of failures encountered during the iterative repair process across various settings. The following elucidates the key findings derived from this investigation.
Figure 4. Error types observed in failed repair iterations across all settings
Figure 4 illustrates the distribution of failed iterations across distinct settings and provides a comprehensive breakdown of failure reasons. Predominantly, failures arise when the proposed repairs fail to fulfill assertions, resulting in the generation of counterexamples. Additionally, syntax errors in proposed repairs impede the compilation process by Alloy Analyzer. Settings 1-3, employing GPT-4-32k, manifest instances of failures attributed to incorrect message formats, a phenomenon absent in Settings 4-6 utilizing GPT-4-Turbo. Such disparities suggest discrepancies in the compliance of LLM responses with the JSON format mandated by the alloy_analyzer_tool. Categories such as ”Repetition” and ”No instances” exhibit lower prevalence. ”Repetition” denotes instances where proposed repairs mimic defective models, while ”No instances” signify scenarios where proposed models fail to generate instances, indicative of an incapacity to meet model constraints.
Figure 5. Repair progress in Setting-6 (GPT-4-Turbo and Auto-Feedback) across iterations. Percentage numbers show transition volumes between statuses in subsequent iterations. Cumulative counts of fixed models follow each ’Fixed’ block.
Figure 5 tracks outcomes across all iterations in Setting-6, revealing the evolving nature of failure categories during the iterative prompting process. While instances of ”Counterexample” failures are effectively addressed, challenges persist in rectifying issues classified under ”Repetition.” Moreover, a discernible transition of cases between distinct failure categories underscores the dynamic nature of failure characteristics throughout the repair endeavor.
Analysis of failure characteristics during the repair process highlights common issues such as proposed repairs failing assertions and syntax errors, with observed transitions between failure categories over iterative prompting.
5.5. Results for RQ5: Repair Costs
To assess the financial implications of utilizing the APR pipeline with different LLMs, we consider two primary cost factors: (i) time taken for bug resolution and (ii) monetary expenses associated with token consumption, based on OpenAI’s pricing as of March 2024.
Time Analysis: We measure the time taken for bug resolution using the results presented in Table 6, which leverages GPT-4 models. Figure 6 illustrates the runtime analysis, contrasting fixed and unfixed scenarios. Across different settings, the median duration required for bug repair varies from 37.3 seconds in Setting-1 to 120.3 seconds in Setting-3. Notably, the maximum time spent on repairs reached 493.28 seconds in Setting-5. Conversely, the median duration associated with unfixed bugs amounted to 273.8 seconds and 392.2 seconds in Settings-1 and 4, respectively.
Figure 6. Distribution of running time across all settings for fixed bugs (left) and unfixed bugs (right).
Monetary Costs: Figure 7 presents the monetary costs incurred under each setting for both fixed and unfixed models. Settings utilizing GPT-4-Turbo show the lowest costs for both fixed and unfixed models, attributed to the cheaper cost per token for GPT-4-Turbo compared to GPT-4-32k. Additionally, settings employing the prompt-agent (Settings-3 and 6) exhibit marginally higher costs than their counterparts utilizing the same LLM.
Figure 7. Monetary distribution across all settings for fixed bugs (left) and unfixed bugs (right).
The analysis of repair costs in RQ5 reveals variations in both time and monetary expenses across different settings, with GPT-4-Turbo exhibiting the lowest costs for bug resolution, and settings utilizing the prompt-agent showing slightly higher expenses.
6. Threats to Validity
Internal Validity. Variations in prompt design and format may introduce biases in the repair process, affecting the outcomes independently of the LLM’s capabilities. To mitigate this threat, we carefully standardized the prompt design across all experiments, ensuring consistency in the information provided to the LLMs. Additionally, we conducted analyses to assess the impact of prompt variations on repair performance, enabling us to isolate the effects of LLM capabilities from potential biases introduced by prompt design.
External Validity. The generalizability of our findings may be limited by the specific characteristics of the benchmarks used. To address this, we utilized diverse benchmark suites and evaluated the performance of the repair pipeline across various scenarios. Furthermore, we documented the experimental setup comprehensively to facilitate replication and external validation of our results.
Construct Validity. Biases in the measurement of repair performance metrics could distort our assessment of LLM capabilities. To mitigate this, we employed standardized metrics and conducted sensitivity analyses to validate the robustness of our findings. Additionally, we ensured transparency and reproducibility in our methodology to enhance the validity of our measurements.
7. Related Work
Several recent studies have explored the integration of Large Language Models (LLMs) into software engineering tasks, particularly in the realm of program repair. Xia et al. (Xia and Zhang, 2023) enhanced automated program repair (APR) performance by incorporating test feedback into prompts, while Kang et al. (Kang et al., 2023) enabled LLMs to utilize a debugger for information gathering and patch generation. Additionally, Fuzz4All (Xia et al., 2024) leverages LLMs as an input generation and mutation engine, employing an auto-prompting phase to generate concise input prompts. RepairAgent (Bouzenia et al., 2024) employs LLMs, agents, and dynamic prompts for APR, albeit in the context of repairing Java applications. TestPilot (Schäfer et al., 2024) utilizes LLMs for generating unit test cases for JavaScript, employing a few-shot learning approach to refine prompts with failing tests and error messages.
In contrast to these efforts, our study focuses specifically on using LLMs for automated specification repair (ASR), a less explored area in software engineering. We adopt an approach similar to Fuzz4All’s auto-prompting phase, albeit with modifications to suit the requirements of repairing Alloy declarative specifications. Moreover, while existing works predominantly use LLMs for program repair in imperative languages like Java and JavaScript, our study extends the application of LLMs to the domain of Alloy specifications, addressing a gap in the literature regarding ASR techniques for declarative languages.
8. Conclusion
In this study, we explore the potential of pre-trained LLMs to facilitate the repair of Alloy specifications, taking into account recent advancements in LLMs, including the use of agents, feedback mechanisms, zero-shot learning capabilities, and auto-prompting techniques. The comparative evaluation highlights the superior performance of the GPT-4 model family over GPT-3.5-Turbo, underscoring their promising applicability. The investigation reveals that employing a multi-agent repair pipeline enhances the repair process, albeit with a marginal increase in token consumption. Overall, the findings of this research indicate a positive outlook for the application of LLMs in automated program repair for declarative specifications, particularly through repair pipelines that integrate contemporary innovations in the field of LLMs.